OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

OpenStudy (potatoes.ramu):

Multiply the second equation throughout by 2

OpenStudy (anonymous):

I don't understand the exponents part... I got x^4 + 10x = 28

OpenStudy (potatoes.ramu):

And then subtract the two equations :)

OpenStudy (anonymous):

OOOOhhhhh!

OpenStudy (anonymous):

Thank you!

OpenStudy (potatoes.ramu):

Wait

OpenStudy (potatoes.ramu):

Is it 2x or x^2? in the first equation?

OpenStudy (anonymous):

oh sorry x^2

OpenStudy (anonymous):

its y^2 + x^2

OpenStudy (potatoes.ramu):

Okay in the second equation

OpenStudy (potatoes.ramu):

Square both sides

OpenStudy (potatoes.ramu):

You need to use the algebraic identity

OpenStudy (anonymous):

y^2 - x^2 = 25

OpenStudy (ranga):

From the second equation you get y = x + 5 Put this in (1): y^2 + x^2 = 53 (x+5)^2 + x^2 = 53 expand, simplify and solve the quadratic equation.

OpenStudy (potatoes.ramu):

i.e, (a-b)^2= a^2+b^2-2ab

OpenStudy (potatoes.ramu):

Then you just subtract one from the other :)

OpenStudy (anonymous):

ok I think I got it now thank you :)

OpenStudy (potatoes.ramu):

Best of luck! :)

hartnn (hartnn):

sonshine, do what ranga said.